OpenStudy (anonymous):

3x - square root of (-25) = 6 -2yi Find x and y i got 3x-5i=6-2yi how can i get x and y from here

OpenStudy (nenadmatematika):

3x-5i=6-2iy so -5i=-2iy and 3x=6 so y=5/2 and x=2

OpenStudy (anonymous):